HI! I need your help to solve this word problem. I have a lot of trouble understanding how to short the information out.
Word Problem:
At a university, 1200 students are enrolled in engineering. There are twice as many in electrical engineering (e) as in mechanical engineering, and three times as many in chemical engineering (c) as in mechanical engineering (m).
QUESTIONS:
1. Which system of equations represents the number of students in each program?
A)c+m+e=1200,2m=e,3m=c
B)c+m+e=1200,3m=e,2m=c
C)c+m+e=1200,2e=m,3c=m
D)c+m+e=1200,2m=e,3m=2e
2. How many students are enrolled in the mechanical engineering program?
A)200
B)400
C)600
D)1200

Thank You for you Time !:)

Found 2 solutions by richwmiller, Alan3354:
Answer by richwmiller(17219)

We all agree that c+m+e=1200
There are twice as many in electrical engineering (e) as in mechanical engineering, and
Does this mean 2e=m or 2m=e?
three times as many in chemical engineering (c) as in mechanical engineering (m).
Does this mean 3c=m or 3m=c?
We can eliminate other choices such as 3m=e in B) and 3m=2e in D)
So now we need to choose between
A)c+m+e=1200,2m=e,3m=c
C)c+m+e=1200,2e=m,3c=m

Does "There are twice as many in electrical engineering (e) as in mechanical engineering" mean there are more e or more m?
There are more e so it would take 2m to equal e or 2m=e
So A) is the answer so far.
three times as many in chemical engineering (c) as in mechanical engineering
Are there more c or more m
There are more c than m
so it would take 3m to equal c or 3m=c
which again is in A)
there are 200 mechanical engineers


Answer by Alan3354(69443)

At a university, 1200 students are enrolled in engineering. There are twice as many in electrical engineering (e) as in mechanical engineering, and three times as many in chemical engineering (c) as in mechanical engineering (m).
QUESTIONS:
1. Which system of eqations represents the number of students in each program?
A)c+m+e=1200,2m=e,3m=c
B)c+m+e=1200,3m=e,2m=c
C)c+m+e=1200,2e=m,3c=m
D)c+m+e=1200,2m=e,3m=2e
-------------------------
There are twice as many in electrical engineering (e) as in mechanical engineering --> e = 2m
------------------------
and three times as many in chemical engineering (c) as in mechanical engineering (m) --> c = 3m
-----------------------------
-----------------------------
2. How many students are enrolled in the mechanical engineering program?
A)200
B)400
C)600
D)1200
----------
c+m+e = 1200
3m + m + 2m = 1200
6m = 1200
Close enough?